OpenStudy (anonymous):

what is the 5th partial sum of sigma notation with upper limit of infinity, lower limit of n=1, of -4+5n

OpenStudy (anonymous):

\[ a_n= 4-5n\\ s_5=a_1+ a_2 + a_3 + a_4 + a_5\\ \] Can you finish it now?

OpenStudy (anonymous):

\[ a_1= 4 - 5(1)= -1\\ a_2= 4 - 5(2)= -6\\ a_3= 4 - 5(31)= -11\\ a_4= 4 - 5(4)= -16\\ a_5= 4 - 5(5)= -21\\ s_5 = -1-6-11-16-21=-55 \]